ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 68, 94 ,03 Sharipbaev, Altynbek, Kazakh state Academy of Control, Kazakhstan A verification of the Refal-programms The verification of the Refal-proogramms is decided by logic languages: for the description of the semantics of the language Refal and for the proof of the conditions of verification of the refal-programs. Exists algorithm which finds out comparability of given input to the left parts of refal-rules. For any X p(X, P (X)) is true if and only if [P, p](X) is executable. The program P is partial correctly if and only if for any X Notf(X)Or[P, g](Y) is executable. The program P is total correctly if and only if for some X f(X)And[P, g](X) is not executable. Reference: V. F. Turchin, Refal-5, Programming Guide and Reference manual, New England publishing: Holyoke, Massachusetts